### v2

# Dial-In LLM: Human-Aligned LLM-in-the-loop Intent Clustering for Customer Service Dialogues

Abstract
Discovering customer intentions in dialogue conversations is crucial for automated service agents. Yet, existing intent clustering methods often fail to align with human perceptions due to the heavy reliance on embedding distance metrics and sentence embeddings. To address these limitations, we propose integrating the semantic understanding capabilities of LLMs into an LLM-in-the-loop (LLM-ITL) intent clustering framework. Specifically, this paper (1) investigates the effectiveness of fine-tuned LLMs in semantic coherence evaluation and intent cluster naming, achieving over 95% accuracy; (2) designs an LLM-ITL clustering algorithm that facilitates the iterative discovery of coherent intent clusters; and (3) proposes task-specific techniques tailored for customer service dialogue intent clustering. Since existing English benchmarks pose limited semantic diversity and intent labels, we introduced a comprehensive Chinese dialogue intent dataset, comprising over 100,000 real customer service calls and 1,507 human-annotated intent clusters. The proposed approaches significantly outperformed LLM-guided baselines, achieving notable improvements in clustering quality and a 12% boost in the downstream intent classification task. Combined with several best practices, our findings highlight the potential of LLM-in-the-loop techniques for scalable and human-aligned problem-solving. Sample code and datasets are available at: anonymous.4open.science/r/Dial-in-LLM.
Dial-In LLM: Human-Aligned LLM-in-the-loop Intent Clustering for Customer Service Dialogues
Mengze Hong1  Di Jiang2thanks: Corresponding author: dijiang@webank.com  Yuanfeng Song2  Lu Wang3  Wailing Ng1 Yanjie Sun1  Chen Jason Zhang1  Qing Li1 1Hong Kong Polytechnic University  2AI Group, WeBank Co., Ltd  3Shenzhen University
1 Introduction
Intent discovery is an essential task in many NLP applications, such as designing AI-driven dialogue systems Hengst et al. (2024) and learning utterance patterns Ghosal et al. (2020). These systems often rely on an intent classification model that assigns intent labels. However, obtaining labeled data for training such supervised models is difficult, especially for dialogue data where topics evolve and contexts change Hengst et al. (2024); Hong et al. (2024). This motivates the need for robust intent discovery methods that can adapt to dynamic dialogue variations and applications.
Figure 1: Comparison of embedding-distance metric and human perception: cosine similarity failed to identify the same intentions under linguistic diversity (top), as well as distinguish different intentions under semantic diversity (bottom).
Intent clustering, which has evolved from unsupervised text clustering, offers significant advantages in automatically discovering thematic relationships Gung et al. (2023). Although text clustering is widely studied, existing research mainly concentrates on developing meaningful representations for text sentences Yin et al. (2021), treating it as a conventional clustering problem without addressing distinctive characteristics such as linguistic and semantic diversity (see Figure 1). Moreover, clustering algorithms heavily rely on quantitative optimization of embedding distance metrics with encoded representations, which is believed to hinder human-interpretable evaluation and validation.
With the growing research interest in integrating Large Language Models (LLMs) into the problem-solving pipeline Zhang et al. (2023); Viswanathan et al. (2024), known as the LLM-in-the-loop (LLM-ITL) paradigm Hong et al. (2025), this paper proposes an LLM-ITL solution specifically designed for dialogue intent clustering. This approach emphasizes the use of intermediate clustering results with human-aligned LLM utilities, providing practical benefits to industrial applications. The main contributions of this work include:
  1. 1.
    We introduce the largest Chinese dialogue intent clustering dataset, notable for its high semantic diversity and a significant inclusion of noisy, out-of-domain intents. This dataset is derived from over 100,000 real-world customer service call sessions, comprising 1,507 intent clusters annotated by human experts. It fills a crucial gap in evaluation benchmarks for complex intent clustering and classification.
  2. 2.
    We demonstrate the effectiveness of fine-tuned LLMs in assessing the semantic coherence of intent clusters and accurately providing intent labels across various sampling strategies. This highlights two effective and efficient utilities for designing LLM-in-the-loop solutions tailored for intent clustering.
  3. 3.
    We propose an LLM-in-the-loop clustering paradigm under both context-free and context-aware settings, which effectively utilizes intermediate results during the clustering process. This method surpasses state-of-the-art baselines and offers significant practical advantages in downstream applications.
2 Related Work
This study is grounded in two key areas of research: LLM-guided text clustering and intent clustering.
LLM-guided Text Clustering.
The discussion of text clustering with LLMs integration emerged as a trending topic since 2023. For instance, Zhang et al. (2023) utilized instruction-tuned LLMs like ChatGPT to better align clustering granularity with user preferences by refining the embedding space, while Viswanathan et al. (2024) demonstrated that LLMs can significantly enhance clustering quality by improving input features, providing constraints during clustering, and performing post-corrections. Feng et al. (2024) proposed a simple approach to refine edge points with LLM, resulting in minor performance improvement. Recently, Hong et al. (2025) proposed the concept of LLM-in-the-loop machine learning, categorizing LLM integration into data-, model-, and task-centric approaches. A “human-in-the-loop” solution can be replicated by LLM utilities Chen et al. (2024), enabling convenient and cost-efficient industrial deployment.
Intent Clustering.
Intent clustering is a fundamental step in intent induction Chandrakala et al. (2024) and the development of intent classifiers Gung et al. (2023). While various techniques have been proposed concerning semi-supervised Kumar et al. (2022) and deep learning approaches Lin et al. (2020); Zhang et al. (2021), recent advancements in LLM-based intent clustering systems have shown significant practical benefits Liang et al. (2024). Notably, the IDAS method proposed by De Raedt et al. (2023) is distinguished by its innovative use of LLMs and in-context learning to generate descriptive utterance labels, thereby enhancing clustering performance. However, the effective use of LLMs continues to pose challenges, such as managing input length and ensuring model robustness Song et al. (2023).
Highlights.
Previous studies have primary focused on a data-centric approach, prioritizing the enhancement of input text representations. In contrast, this paper emphasizes on the importance of intermediate results during the clustering process, designing tailored solutions for human-aligned intent clustering. Additionally, rather than relying on public benchmark datasets with limited complexity, this work tackles a more realistic setting by clustering a noisy dataset transcribed from real customer service calls, offering practical solutions in handling dialogue intent clustering.
3 Preliminaries
3.1 Coherence Evaluation
Inspired by human behaviors in perceiving texts and making comparisons based on semantic meaning rather than word-level similarity Peter W. Foltz and Landauer (1998), this paper proposes the semantic consistency within a cluster, known as semantic coherence, as a superior metric and optimization target. The evaluation is simplified by formulating it as a binary classification problem, leading toward intuitive evaluation and easing the training of the LLM evaluator. Good clusters consist of sentences focused on a specific topic, while Bad clusters contain inconsistent or ambiguous intentions (see Table 6 for examples). This task is particularly challenging for traditional machine learning classifiers due to a lack of semantic understanding Mimno et al. (2011), necessitating fine-tuned LLMs Gu et al. (2025) for robust evaluation.
3.2 Cluster Naming
Giving each cluster a concise and meaningful name is essential for many downstream applications Pattnaik et al. (2024). In this paper, we introduce a novel naming convention, "Action-Objective," which is particularly effective for capturing dialogue intents that are typically topic-oriented (e.g., insurance, loan) and involve distinct actions (e.g., inquire). Examples of human annotations are shown in Table 7, and a comparison of different naming conventions is presented in Table 8.
4 Proposed Methods
4.1 Context-Free Intent Clustering
Context-free intent clustering represents the simplest setting, which ignores task-related information such as dialogue roles and domain-specific knowledge. The performance of the context-free approach provides a general expectation of effectiveness when applying the proposed methods to different datasets and tasks.
4.1.1 LLM-in-the-loop Intent Clustering
With the introduction of cluster-level semantic coherence evaluation, individual clusters that are identified to be badly clustered can be further refined, enabling the iterative discovery of high-quality clusters. The proposed method is presented in Algorithm 1. To determine the appropriate number of clusters, which is often unknown in real-world applications, a candidate space of cluster numbers $\bm{N}$ is predefined. In each iteration, text embeddings $\bm{t}$ are input into a clustering algorithm $\bm{F}(\bm{t},n_{i})=\bm{c_{n_{i}}}$ for each candidate cluster number $n_{i}$. The coherence of the clustering results is then evaluated by a fine-tuned LLM, and the cluster assignment yielding the highest number of good clusters is selected as the intermediate result. The good clusters in this result are retained, and the sentences in the bad clusters are re-clustered in the subsequent iteration, ensuring that the final result consists mostly of high-quality clusters. The final clusters are labeled with LLM naming to improve the interpretability and practical values.
Determining the optimal number of clusters in a noisy text corpus is often challenging. Here, we propose a local search heuristic that maximizes the “good/bad” ratio at each iteration:
$\max_{n_{k}\in\bm{N}}\quad\frac{\#\,\bm{c_{n_{k}}}|\,g_{k}=good}{\#\,\bm{c_{n_{k}}}|g_{k}=bad}$
where $n_{k}$ denotes the best cluster number at the $k$th iteration. This enables the automatic discovery of suitable cluster numbers in a step-by-step manner. The search space $\bm{N}$ should chosen carefully to balance accuracy and efficiency; thus, we propose a search space pruning strategy, as illustrated in Section 6.4.2 for more details.
Algorithm 1 LLM-in-the-loop Intent Clustering
1:sentences, $LLM_{eval}$, $LLM_{name}$, cluster numbers $\bm{N}=[n_{1},\dots,n_{k}]$
2:set of clusters $\bm{C}$ with intent labels $\bm{L}$
3:$\bm{t}\leftarrow\text{SentenceTransformer}(sentences)$
4:$\bm{C}\leftarrow\emptyset$
5:for each $n_{i}\in N$ do
6:     initial clusters $\bm{c_{n_{i}}}\leftarrow\bm{F}(\bm{t},n_{i})$
7:     $\bm{g_{i}}\leftarrow[LLM_{eval}(c_{1}),\dots,LLM_{eval}(c_{n_{i}})]$
8:end for
9:$i\leftarrow\mathop{\arg\max}_{i}\frac{\#\,\bm{c_{n_{i}}}|\,g_{i}=good}{\#\,\bm{c_{n_{i}}}|g_{i}=bad}$
10:$\bm{C}\leftarrow\bm{C}\oplus\bm{c_{n_{i}}}\;|\,g_{i}=good$
11:if #$\#\,\bm{c_{n_{i}}}|\,g_{i}=bad>\epsilon$ then
12:     reverse to 3 with $\bm{t}\leftarrow\bm{c_{n_{i}}}\;|\,g_{i}=bad$
13:end if
[…]
4.1.2 Post-Correction with Intent Labels
From an initial exploration, multiple clusters representing similar intents may emerge in the final result (see Table 9). This occurs because, as the iterative process progresses, the number of poorly clustered sentences decreases, resulting in smaller and less representative clusters being formed. The challenge of embedding distance prevents the grouping of these clusters, motivating post-correction by merging similar clusters based on intent labels.
Unlike Viswanathan et al. (2024) and Feng et al. (2024), which directly query LLMs the correctness of cluster assignments (i.e., whether the cluster assignment of a data point is correct), we leverage the cluster naming capability of LLMs to label each cluster with its underlying intent. These labels $L(c_{i})$ can be merged when $\text{Dist}(L(c_{i}),L(c_{j}))<\theta$, where $\theta$ is a predefined merge threshold and Dist is a distance metric. As the intent label is a concrete and concise summary of the whole cluster, the embeddings of these labels are believed to retain semantic information and are suitable for merging similar clusters.
4.2 Context-Aware Intent Clustering
From a practical perspective, dialogue datasets are often accompanied by domain-specific features or use-case scenarios (e.g., customer service calls, group discussions). This motivates the incorporation of contextual information into the clustering process. However, the discovery of suitable knowledge remains a challenging task; thus, a knowledge discovery process is proposed based on LLM naming, where intermediate clustering assignments are utilized to achieve improved clustering results.
The customer service dialogues typically involve only the customer and the service agent Lin et al. (2022). The classification task that assigns a sentence to its associated role is relatively simple with labeled training data. To provide an unsupervised solution, the sentence roles are determined based on the labeled intent clusters. By heuristics, sentences within clusters labeled as “inquire-” or similar actions are mostly sent from the customer, and vice versa. A two-step approach is proposed: initially, intermediate clustering results are obtained using K-means or a context-free approach, and these results are divided based on intent labels, forming two clusters with distinct roles. In the second step, the sentences corresponding to each role are clustered again to achieve more refined outcomes, allowing for a clear separation of customer intent from service agent intent.
5 Experiment
5.1 Dataset
The proposed dataset contains 1,507 high-quality intent clusters manually annotated from over 100,000 customer service calls, comprising 55,085 distinct sentences with an average length of 17 Chinese characters per sentence (see Appendix A for annotation details) [^1] The full dataset will be open-sourced upon acceptance. Sample dataset can be found at GitHub.. Among these intents, 885 are identified as domain-specific queries (e.g., “inquire-insurance”), while 622 are categorized as out-of-domain (e.g., “provide-location”), representing general queries that might commonly arise in a customer service interaction. This dataset is the first Chinese benchmark for intent clustering and the largest of its kind compared to the existing English benchmark datasets shown in Table 1, which presents several new challenges such as lexical sparsity in short-text clustering, dependence on prior knowledge for context enrichment, and balancing accuracy with computational efficiency. The semantic diversity, calculated based on the average cosine distance between individual sentences and the cluster centroid Casanueva et al. (2022), depicts the complexity of this dataset.
Number of Number of Semantic
Dataset examples intents diversity
BANKING77 13,083 77 0.209
OOS 23,700 150 0.254
NLU++ 3,080 62 0.367
ours 55,085 1507 0.538
Table 1: Comparison of our proposed dataset with existing benchmark intent clustering datasets.
5.2 Metrics
Normalized Mutual Information (NMI) measures the degree of similarity between ground-truth and predicted clusters, scaled between 0 (no mutual information) and 1 (perfect correlation). However, researchers have claimed biased behavior in NMI Jerdee et al. (2024), necessitating further evaluation that reflects the true quality of the intent cluster from human perception.
Goodness score measures the proportion of good clusters evaluated by the fine-tuned LLM evaluator. For intermediate steps of iteration where the number of clusters is unknown, the good/bad ratio is used as an invariant measure of clustering quality:
$\text{goodness}_{i}=\frac{\text{\# good clusters}}{\text{\# bad clusters}}$
For the final clustering results, the percentage of good clusters among all clusters is reported. Note that the evaluation uses a different LLM evaluator than the one used for intermediate evaluation to ensure the fairness of the reported metric:
$\text{goodness}_{final}=\frac{\text{\# good clusters}}{\text{\# total clusters}}$
This metric offers a comprehensive understanding of the quality of the produced intent clusters, allowing for accurate cluster-level evaluation. Additionally, it can be easily deployed in applications without ground-truth annotations, thereby eliminating the need for human involvement.
Category Method NMI #Good
Base models K-Means 0.7899 94.8%
GMMs 0.7903 91.1%
Hierarchical* 0.8001 94.9%
Data-centric ClusterLLM Zhang et al. (2023) 0.7284 91.2%
IDAS De Raedt et al. (2023) 0.8109 93.6%
Keyphrase Viswanathan et al. (2024) 0.8371 94.5%
Context-Free LLM-ITL (random) 0.8202 97.7%
LLM-ITL (convex) 0.8208 97.8%
LLM-ITL + keyphrase 0.8378 96.4%
Context-Aware LLM-ITL + role 0.8679 97.2%
LLM-ITL + role + merge 0.8826 97.6%
Table 2: Performance of baselines and proposed methods in dialogue intent clustering. The best results are bolded.
6 Results and Discussions
6.1 LLM Utilities
{CJK}
UTF8gkai
LLM qwen7b qwen14b baichuan2-7 chatglm3-6b
Accuracy 96.25% 97.50% 89.17% 95.83%
Table 3: Performance of fine-tuned LLMs in evaluating semantic coherence of intent clusters.
Coherence Evaluation.
To enhance evaluation performance while preserving data privacy, four Chinese open-sourced LLMs Liu et al. (2024) are fine-tuned using LoRA Hu et al. (2022) on a training dataset comprising 1,772 intent clusters annotated as either “good” or “bad”. Table 3 presents performance of fine-tuned LLMs based on 480 unseen clusters. The results suggest that mainstream open-source LLMs can serve as robust evaluators for assessing the quality of intent clusters.
{CJK}
UTF8gkai
LLM qwen7b qwen14b baichuan2-7 chatglm3-6b
Accuracy 92.8% 94.3% 94.3% 94.4%
Table 4: Performance of fine-tuned LLMs in naming intent clusters.
Cluster Naming.
Similar to coherence evaluation, we manually annotated 2,500 clusters with corresponding intent labels to fine-tune and assess the utility of LLMs for cluster naming. Since labels are not unique, the accuracy is manually assessed by four human experts based on the alignment with the true labels in the dataset. The results in Table 4 suggest that fine-tuned LLMs have promising performance in providing intuitive names and aligning closely with human perceptions.
6.2 Intent Clustering Performance
Implementation.
The qwen14B model was utilized for both coherence evaluation and cluster naming. For all clustering methods, semantic-rich embeddings are obtained using the BAAI General Embeddings (BGE) model [^2] https://huggingface.co/BAAI/bge-large-zh-v1.5 Xiao et al. (2024). Among the clustering algorithms, hierarchical clustering demonstrated the highest NMI and goodness scores, thus serving as the base model for the proposed methods. Sampling techniques (i.e., random and convex) are implemented to select representative sentences from each cluster as input for LLM.
Among data-centric baselines, refining sentence embeddings through keyphrase expansion demonstrated the most improvement. However, deriving pairwise constraints for fine-tuning the embedding model, as proposed in ClusterLLM, proved ineffective due to the excessive number of clusters in the dataset. Interestingly, the evaluation of goodness does not always align with the NMI score, offering additional insights into cluster quality. For example, while keyphrase expansion improved the NMI score, the generated keyphrases could distort the meaning of the original sentence, resulting in less representative clusters and a slightly lower goodness score compared to the base model.
The proposed LLM-in-the-loop intent clustering, abbreviated as the “LLM-ITL” approach, demonstrated satisfactory improvement in NMI and significant enhancement in the goodness score. In a context-free setting, the integration of LLM evaluation and naming outperformed both ClusterLLM and IDAS baselines. By incorporating effective keyphrase expansion, the NMI improved further, highlighting the potential of integrating data-centric methods into the pipeline. Notably, the proposed method does not require prior knowledge of the number of clusters and conducts parameter searches during the clustering process, whereas the baseline methods preset the number of clusters at $n=1507$. This adaptability underscores its significant application value for extracting intent clusters from any noisy text corpora Akama et al. (2020).
Furthermore, context-aware approaches were evaluated with the goal of enhancing the task performance of customer service dialogue intent clustering. A consistent improvement was observed when incorporating dialogue role into the clustering process, resulting in a 4.48% increase in NMI without compromising the goodness score. This highlights the importance of task-centric design in effectively addressing industrial problems.
6.3 Application Performance
One objective of intent clustering is to create a high-quality labeled dataset for training intent classifiers that can effectively handle future data inputs Gung et al. (2023). To demonstrate the practical advantages of the proposed methods, we further evaluate their performance by training BERT classifiers (bert-base-chinese model) based on the clustered results. Compared to the training data produced by the baseline method, which achieved only 65% classification accuracy, the proposed LLM-in-the-loop method improved the accuracy by 12%, providing promising evidence of the improved cluster quality.
6.4 Ablation Studies
We present two ablation studies to derive best practices for industrial deployment using different sampling and hyperparameter tuning strategies.
6.4.1 Effective LLM Evaluation
Optimizing the input to LLMs is crucial for enhancing performance and reducing costs Song et al. (2025); Jiang et al. (2024). As shown in Figure 2, convex sampling outperforms random sampling in selecting representative sentences for coherence evaluation. Interestingly, as the sample time $t$ (for repeated evaluation) or the dimension of convex hull $d$ increases, performance consistently declines. This suggests that simpler, more efficient parameters lead to better accuracy. This pattern is verified through experiments with the ChatGLM model (see red line), which exhibits the same pattern.
To better assess the consistency and alignment between LLM judgments and human perceptions, we utilize the concept of crowdsourcing for collaborative coherence evaluation Zhang et al. (2025). The proposed LLM crowd consists of diverse entities Zhang et al. (2024) represented by various fine-tuned LLMs and uses a majority voting mechanism to aggregate their judgments Schoenegger et al. (2024). Based on the evaluation, we argue that a single fine-tuned LLM can effectively serve as a robust evaluator, as evidenced by the comparable performance to crowdsourced accuracy (cross markers in Figure 2) when using convex sampling.
6.4.2 Parameter Search Space Pruning
In practice, the search space for the number of clusters can be extensive, ranging from a few options to several hundred. To mitigate the computational costs associated with traditional model-based selection, we propose an LLM-native method for search space pruning, which predicts the optimal cluster number for subsequent iterations using logs from previous iterations (see Appendix 10). By pruning the search space and selecting the top 5 non-repetitive predicted solutions, we achieved an accuracy of 89.47% (see Table 5), indicating a high likelihood that the optimal cluster number is among these candidates. This approach substantially reduces redundant model fittings and enhances the efficiency of deploying the proposed methods.
{CJK}
UTF8gkai
Top1 Top2 Top3 Top4 Top5
Accuracy 26.32% 47.37% 73.68% 84.21% 89.47%
Table 5: Search space pruning by LLMs.
7 Conclusion
This paper tackles the dialogue intent clustering task by leveraging intermediate results through a human-aligned LLM-in-the-loop approach. Extensive experiments show that fine-tuned LLMs perform robustly in semantic coherence evaluations and cluster labeling. Both context-free and context-aware methods surpass existing LLM-guided clustering baselines, significantly improving cluster quality and downstream application performance. The ablation studies further provide practical insights for efficient industrial deployment. Future research is encouraged to enhance LLM reliability by integrating confidence estimation into the evaluation process, as well as employing advanced base algorithms and more sophisticated judgment formats to reveal additional insights.
Appendix A Data Collection and Annotation
This section details the three-stage data collection and annotation process for the proposed Chinese dialogue intent clustering dataset.
Data Collection.
The original dialogue dataset is sourced from audio transcriptions of customer service calls across several domains. Initially, the dataset contains over 120,000 calls. Strict filtering is applied to maintain data integrity and exclude confidential or sensitive content, resulting in 8,184 dialogues and 69,839 sentences. Subsequent data cleaning removes duplicate sentences, resulting in a final set of 55,085 unique sentences.
Data Annotation.
To create the intent clustering dataset, we recruited 15 human experts, each with over five years of experience in customer service call centers, to participate in the data annotation process:
  1. 1.
    K-means clustering was initially applied with $n=2000$, serving as a starting point for the data annotation.
  2. 2.
    Each expert assessed the initial clusters by inspecting the semantic coherence of the sentences. They were instructed: “Label the cluster as ’Good’ if all sentences share the same intent; otherwise, label it as ’Bad’. This resulted in 1283 good clusters and 717 bad clusters.
  3. 3.
    For the good clusters, annotators were asked to label the underlying intent using the naming convention “Action-Objective.” Five samples were provided as guides. After intent annotation, 1255 clusters with unique intentions were retained, and 28 clusters were merged due to replicated intent.
  4. 4.
    For the bad clusters, annotators reassigned each sentence to the appropriate intent cluster based on the annotated labels. Sentences that did not fit into any preexisting clusters were assigned to new clusters, and the same process was repeated as in Step 2.
Data Verification.
In the final stage, a separate group of 10 experts reviewed the annotated clusters for accuracy and consistency. Any discrepancies were resolved through consensus, ensuring the dataset’s reliability and validity for further analysis. The final dataset consists of 1,507 high-quality intent clusters.
Appendix B Prompt Template and More Results for LLM Utilities
The prompt for coherence evaluation and cluster naming are translated into English for demonstration. For each input, a few-shot demonstration with five input-output pairs is provided to ensure the consistency of the output format and enhance understanding of the task.
Coherence Evaluation -  Your are a helpful assistant for sentence clustering. Based on the relevancy and common points of the following sentences in a cluster, classify the cluster as: “Good” or “Bad”. Only provide the label without any additional content. Example: input:[sentences] output:[label] input:{[sentences]} output:
Cluster Naming -  Your are a helpful assistant for sentence clustering. Based on the relevancy and common points of the following sentences in a cluster, summarize the cluster with an “Action-Objective” label. Only provide the label without any additional content. Example: input:[sentences] output:[label] input:{[sentences]} output:
Considering the resource constraints of industrial deployment, we evaluated the coherence performance using the OpenAI APIs: GPT-3.5, GPT-4, and GPT-4o, achieving accuracies of 89.58%, 93.54%, and 94.17%, respectively. These findings indicate that a smaller, fine-tuned LLM, such as qwen7b, can outperform these advanced models and potentially reduce the costs associated with API requests [^3] The fine-tuned LLMs for coherence evaluation and cluster naming will be released upon acceptance.. To highlight the importance of fine-tuning, we also tested a vanilla Qwen2.5 model without fine-tuning, which resulted in a significantly lower accuracy of 75.63%.
Model Sampling Method 1 2 3 4 5
qwen14b convex - 96.25% 95.42% 95.42% 95.21%
random (n=10) 95.83% 94.38% 93.33% 92.92% 92.92%
random (n=20) 94.58% 95.00% 94.38% 94.17% 94.17%
chatglm3-6b convex - 92.08% 90.83% 91.04% 90.42%
random (n=10) 92.29% 88.75% 85.83% 84.58% 85.21%
random (n=20) 90.42% 89.58% 89.38% 89.58% 88.98%
Figure 2: Performance comparison of sampling methods and hyperparameters for effective LLM evaluation.
{CJK}
UTF8gkai
Cluster Coherence Original Sentences English Translation
Good "给企业固定资产买保险,大约能投大约得投保多少呢", "就是假如我有一百万的企业固定资金买保险大概投保多少钱啊", "请问一下我想咨>询一下企业财产保险", "您好我想了解一下这个企业财产保险", "就了解一下这个企业财产保险","你们是财险", "是哦这个属于财险了对吧", "企业财产保险的", "这些都属于财产险对吗", "财险人工那你这是", 就是如果我要为我这个私营企业买这个保险需要什么手续", "财产险", "呃企业财产保险是以什么为保" "If I buy insurance for a company’s fixed assets, how much insurance will it cost?", "That is, if I have a million corporate fixed assets, how much will it cost to buy insurance?", "Excuse me, I would like to consult> Ask about corporate property insurance", "Hello, I want to know about this corporate property insurance", "Just want to know about this corporate property insurance", "You are a property insurance company", "Yes, this belongs to property insurance "Right?", "Enterprise property insurance", "These all belong to property insurance, right?", "What about property insurance workers?", What do I need if I want to buy this insurance for my private enterprise? Procedure","Property Insurance", "Well, what does corporate property insurance cover?"
Bad "就是连连续,就是一直一直保", "但是它连不上是怎么回事啊?", "哦就是主要是直接给公司转账对吧?", "接吗", "接也是吗?", "我直>接去", "你是直接直接用那个" , "就是从哪接过来再接回去","还是需要从哪儿连线这个宽带", "你直接给我说这些啊", "直接把", "哦直接", "嗯那个礼品是直接就发放了", "是你直接给我回复对吗", "直接到那里去", "嗯,最好直飞。", "请帮我连接+", "把钱打过去的话,我是直接打到那个证券公司", "直接就是" "It’s continuous, it’s always guaranteed", "But what’s wrong with it not being able to connect?", "Oh, it’s mainly to transfer money directly to the company, right?", "Yes", "Yes too?" ?", "I’ll go directly", "Are you using that directly" , "Just connect it from where you are and then connect it back", "Or do you need to connect to the broadband from somewhere", "You Just tell me this directly", "Just give it directly", "Oh directly", "Well, the gift was given out directly", "You replied to me directly, right?","Go there directly" , "Well, it’s best to fly directly.", "Please help me connect +", "If I call the money, I will call the securities company directly", "Directly"
Table 6: Example of “Good” and “Bad” intent clusters in coherence evaluation.
{CJK}
UTF8gkai
Cluster Name Original Sentences English Translation
询问-优惠 (Inquire-Promotion) ’那有什么优惠券什么之类的吗?’, ’是是怎么形式是优惠吗?’, ’还是不是优惠活动,是那个直接给我打我卡里吗 ,还是是什么优惠券儿啊?’, ’就比如新用户他有什么优惠券儿之类的吧!’, ’嗯你你那还有什么优惠的活动吗? 就是比较合适就是合适的动。’ ’Are there any coupons or anything like that? ’, ’What is the form of the discount? ’, ’Still, it’s not a promotion. Is it the one that directly charges my card? , or is it some kind of coupon? ’, ’For example, what coupons does a new user have? ’, ’Well, do you have any other discounts? It is more appropriate and appropriate to move. ’
解答-金额 (Answer-Amount) ’一共是三十一块二毛’, ’就每个月一百三十八’, ’对,一个月也就是四百百四五百块钱嘛, 给您自己做个积累。’, ’十二月份的话是用了三十三块九毛二’, ’对一个月一百三十八’ ’The total is thirty-one and twenty cents’, ’That’s one hundred and thirty-eight cents per month’, ’Yes, that’s four hundred, four hundred and five hundred yuan per month. Make an accumulation for yourself. ’, ’In December, it cost thirty-three dollars and ninety-two cents’, ’That’s one hundred and thirty-eight dollars a month’
Table 7: Example of cluster naming with "Action-Objective" style.
Dataset Naming Convention Example (sentence) Example (label)
NLU scenario-intent Send an email to Alex and write thank you. email sendemail
NLU++ list of keywords How long does it usually take to get a new pin? ["how_long","pin", "arrival","new"]
OOS objective Please tell me why my card was declined yesterday. card_declined
ours action-objective Well, do you have any other discounts? inquire-promotion
Table 8: Comparison of cluster naming convention in intent clustering datasets.
{CJK}
UTF8gkai
Cluster Name Clusters with Similar Intention English Translation
询问-意外事故 (Inquire-Accident) "假如被车碰了或者是被楼上的砖砸了一下", "给别人儿撞的意外", "啊撞到别人然后就是", "就平时有时候开车嘛可能会遇到这个", "嗯哦这种情况,那要是就是我自己不小心撞到了那个某个地方然后", "就是把其他的东西撞到了呀什么的", "那人生意,不是就是,如果是不小心在马上被车撞了的话", "撞到人了吧", "然后不小在行驶当中被别人损害就是说拿石头砸的呀然后", "什么被车撞了之类的,是吗", "被被被撞了,还是被什么一些什么意外事故了", "我把别人的车撞了是吧" "If I were hit by a car or hit by a brick from upstairs," "Accidentally hit by someone else," "Ah, hit someone else and then," "Just sometimes when driving, you might encounter this," "Well, if I’m not careful and hit some place myself," "Just hit something else or something," "That’s a human affair, not just, if it’s accidentally hit by a car on the road," "Hit someone, right?" "Then not small in the process of driving, being damaged by someone else, say, hit with a stone, and then," "What, hit by a car or something, right?" "I hit someone else’s car, right?"
询问-意外死亡 (Inquire-Accident Death) "哦猝死,那猝死算意外吗", "啊,那我想知道那个猝死的话,算是意外死亡吗", "算意外死亡吗", "那如果猝死算是意外死亡吗?", "猝死也算意外事吧", "那那猝死是意外死亡吗如果是猝死的话是", "猝死算是意外死亡吗", "猝死算意外死亡吗", "嗯,那个猝死,猝死属于意外意外险吗,意外死亡吗", "那猝死的话,算意外死亡吗?", "那那那那个就是那个猝死算是意外死亡吗" "Oh sudden death, is sudden death considered accidental?" "Ah, I want to know if sudden death is considered accidental death?" "Is it considered accidental death?" "If sudden death is considered accidental death?" "Sudden death is also an accident, right?" "Is sudden death considered accidental death if it is sudden death?" "Is sudden death considered accidental death?" "Is sudden death considered accidental death?" "Well, that sudden death, does sudden death fall under accidental insurance, accidental death?" "Is sudden death considered accidental death?" "Is that, that, that sudden death considered accidental death?"
Table 9: Example of two high-quality intent clusters with similar intentions.
Epoch 1th 2th 3th
n_cluster good bad rate n_cluster good bad rate n_cluster good bad rate
20 20548 34537 0.595 20 0 2901 0.0 20 9 438 0.021
50 24100 30985 0.778 50 372 2529 0.147 50 70 377 0.186
100 32292 22793 1.417 100 804 2097 0.383 100 79 368 0.215
Best 1600 52184 2901 17.988 800 2454 447 5.490 200 86 361 0.238
Table 10: Example of log record for iterative intent clustering.