### v3

# Dial-In LLM: Human-Aligned LLM-in-the-loop Intent Clustering for Customer Service Dialogues

Abstract
Discovering customer intentions in dialogue conversations is crucial for automated service agents. However, existing intent clustering methods often fail to align with human perceptions due to a heavy reliance on embedding distance metrics and a tendency to overlook underlying semantic structures. This paper proposes an LLM-in-the-loop (LLM-ITL) intent clustering framework, integrating the semantic understanding capabilities of LLMs into conventional clustering algorithms. Specifically, this paper (1) investigates the effectiveness of fine-tuned LLMs in semantic coherence evaluation and intent cluster naming, achieving over 95% accuracy aligned with human judgments; (2) designs an LLM-ITL framework that facilitates the iterative discovery of coherent intent clusters and the optimal number of clusters; and (3) proposes context-aware techniques tailored for customer service dialogue. As existing English benchmarks offer limited semantic diversity and intent groups, we introduce a comprehensive Chinese dialogue intent dataset, comprising over 100k real customer service calls and 1,507 human-annotated intent clusters. The proposed approaches significantly outperform LLM-guided baselines, achieving notable enhancements in clustering quality and lower computational cost. Combined with several best practices, our findings highlight the potential of LLM-in-the-loop techniques for scalable and human-aligned intent clustering.
Dial-In LLM: Human-Aligned LLM-in-the-loop Intent Clustering for Customer Service Dialogues
Mengze Hong1, Wailing Ng1, Chen Jason Zhang1 Yuanfeng Song2, Di Jiang2thanks: Corresponding author: dijiang@webank.com 1Hong Kong Polytechnic University 2AI Group, WeBank Co., Ltd
1 Introduction
Figure 1: Comparison of embedding-distance metric and human perception: cosine similarity failed to identify the same intentions under diverse expression (top), and to distinguish distinct intentions under similar expressions (bottom).
Intent discovery is a crucial task in numerous NLP applications, including the design of AI-driven dialogue systems Hengst et al. (2024) and the learning of utterance patterns Ghosal et al. (2020). While intent clustering techniques are extensively studied to automatically identify thematic relationships within text corpora Gung et al. (2023), existing research primarily focuses on developing meaningful sentence representations and relies on embedding distance metrics for optimization Yin et al. (2021). This approach often overlooks the distinctive characteristics of textual information, such as linguistic patterns and semantic diversity (see Figure 1), and restricts human-aligned evaluation and validation of clustering quality Vinh et al. (2009). This issue is particularly pronounced in languages with rich semantics, such as Chinese, where two seemingly similar sentences can convey entirely different meanings, and vice versa Chen (1993).
With growing research interest in integrating Large Language Models (LLMs) into the problem-solving pipeline Hong et al. (2025), LLM-guided clustering techniques have emerged, demonstrating superior performance over traditional machine learning algorithms. While these methods effectively incorporate language understanding into the clustering process, they primarily focus on data preprocessing, querying LLMs for embedding refinement or data augmentation Zhang et al. (2023); Viswanathan et al. (2024). This approach represents a surface-level integration, potentially missing the benefits of LLMs in providing semantic-driven guidance to the clustering process and the quality of the final results.
In this paper, we introduce an LLM-in-the-loop (LLM-ITL) intent clustering method, designed to facilitate the iterative discovery of coherent intent clusters. Our approach adeptly utilizes intermediate clustering results by incorporating human-aligned LLM utilities, offering computationally efficient and human-interpretable intent clustering. The key contributions of this work are:
  1. 1.
    We introduce the largest Chinese dialogue intent clustering dataset, notable for its high semantic diversity and a significant inclusion of noisy, out-of-domain intents. This dataset is derived from over 100,000 real-world customer service call sessions, comprising 1,507 intent clusters annotated by human experts. It fills a crucial gap in evaluation benchmarks for complex intent clustering and classification.
  2. 2.
    We demonstrate the effectiveness of fine-tuned LLMs in assessing the semantic coherence of intent clusters and accurately providing intent labels across various sampling strategies. This highlights two effective and efficient utilities for designing LLM-in-the-loop solutions tailored for intent clustering.
  3. 3.
    We propose an LLM-in-the-loop clustering algorithm to derive an optimal set of clusters in scenarios where the number of clusters is unknown. This approach effectively leverages intermediate results generated during the clustering process, outperforming state-of-the-art baselines and providing practical benefits for downstream applications. Additionally, the discussion on data sampling and LLM-based crowdsourcing further improves performance.
2 Related Work
This study is grounded in two key areas of research: LLM-guided text clustering and intent clustering.
LLM-guided Text Clustering.
The integration of LLMs into text clustering has become increasingly prominent since 2023. Zhang et al. (2023) introduced ClusterLLM, an innovative approach that leverages instruction-tuned LLMs like ChatGPT to refine sentence embedding spaces through pairwise preference questions, aligning clustering granularity with user preferences. Viswanathan et al. (2024) highlighted the enhancement of clustering quality by LLMs through feature improvement, the imposition of constraints during clustering, and post-correction processes. Additionally, Feng et al. (2024) proposed refining edge points with LLMs, which led to notable performance gains. Recently, Hong et al. (2025) conceptualized LLM-in-the-loop machine learning, classifying LLM integration into data-, model-, and task-centric strategies. This method mirrors a “human-in-the-loop” approach Chen et al. (2024), facilitating effective LLM integration in machine learning applications.
Intent Clustering.
Intent clustering is a fundamental step in intent induction Chandrakala et al. (2024) and the development of intent classifiers Gung et al. (2023). While semi-supervised Kumar et al. (2022) and deep learning methods Lin et al. (2020); Zhang et al. (2021) have significantly advanced intent clustering, recent progress in LLM-based systems has highlighted their potential practical advantages Liang et al. (2024). The IDAS method effectively employs in-context learning to generate descriptive utterance labels and enhance sentence embeddings De Raedt et al. (2023). However, challenges such as computational cost and robustness remain unresolved Song et al. (2023).
Highlights.
Previous studies have primarily focused on enhancing input text representations and the direct use of LLMs for cluster modification, imposing high computational cost and uncertainty. In contrast, this paper implements human-aligned LLM utilities to create a LLM-in-the-loop intent clustering process with emphasis on the intermediate results. Additionally, this work tackles a more realistic scenario by releasing a complex dataset transcribed from real customer service calls, offering practical insights and contributing to more meaningful evaluations in future intent clustering.
Figure 2: Overview of proposed LLM-in-the-loop Dialogue Intent Clustering.
3 Proposed Methods
In this section, we first outline the design of human-aligned LLM utilities. Then, we introduce a comprehensive LLM-in-the-loop intent clustering framework (see Figure 2) and show how each LLM utility contributes to the clustering process.
3.1 LLM Utilities
Coherence Evaluation.
Inspired by human behaviors in perceiving texts and making comparisons based on semantic meaning rather than word-level similarity Peter W. Foltz and Landauer (1998), this paper proposes the semantic consistency within a cluster, known as semantic coherence, as a superior metric and optimization target. The coherence evaluation is formulated as a binary classification problem, leading toward intuitive evaluation and easing the training of the LLM evaluator (see Section 5.6 for different formulations). Good clusters consist of sentences focused on a specific topic, while Bad clusters contain inconsistent or ambiguous intentions (see Table 9 for examples). This task is particularly challenging for traditional machine learning classifiers due to a lack of semantic understanding Mimno et al. (2011), necessitating fine-tuned LLMs Gu et al. (2025) for robust evaluation (see Section 5.1 for fine-tuning details).
Cluster Naming.
Giving each cluster a concise and meaningful name is essential for many downstream applications Pattnaik et al. (2024). In this paper, we introduce a novel naming convention, "Action-Objective," which is particularly effective for capturing dialogue intents that are typically topic-oriented (e.g., insurance, loan) and involve distinct actions (e.g., inquire, confirm). Examples of human annotations are shown in Table 10, and a comparison of different naming conventions is presented in Table 11, demonstrating the superiority of the proposed convention in capturing customer intention with clear goal-oriented expression.
3.2 LLM-in-the-loop Iterative Intent Clustering with Coherence Evaluation
Algorithm 1 LLM-in-the-loop Intent Clustering
1:Unlabeled sentence corpus $\mathcal{S}$, embedding function $f_{\text{emb}}$, coherence evaluator $\mathcal{M}_{\text{eval}}$, candidate cluster numbers $N=\{n_{1},\dots,n_{k}\}$, threshold $\epsilon>0$, max iterations $T_{\text{max}}$
2:Set of clusters $\mathcal{C}$
3:$\mathcal{E}\leftarrow\{f_{\text{emb}}(s)\mid s\in\mathcal{S}\}$
4:$\mathcal{S}^{(0)}\leftarrow\mathcal{S}$, $\mathcal{C}\leftarrow\emptyset$, $t\leftarrow 0$
5:while $\frac{|\mathcal{S}^{(t)}|}{|\mathcal{S}|}>\epsilon$ and $t<T_{\text{max}}$ do
6:     $\mathcal{E}^{(t)}\leftarrow\{f_{\text{emb}}(s)\mid s\in\mathcal{S}^{(t)}\}$
7:     for each $n_{i}\in N$ do
8:         $\mathcal{C}_{n_{i}}^{(t)}\leftarrow F(\mathcal{E}^{(t)},n_{i})$
9:         $\mathbf{g}_{n_{i}}^{(t)}\leftarrow[\mathcal{M}_{\text{eval}}(C_{1}^{(t)}),\dots,\mathcal{M}_{\text{eval}}(C_{n_{i}}^{(t)})]$
10:     end for
11:     $n^{*}\leftarrow\arg\max_{n_{i}\in N}\frac{\sum_{j=1}^{n_{i}}\mathbb{I}\left(\mathbf{g}_{n_{i}}^{(t)}[j]=1\right)}{\sum_{j=1}^{n_{i}}\mathbb{I}\left(\mathbf{g}_{n_{i}}^{(t)}[j]=0\right)+1}$
12:     $\mathcal{C}_{\text{good}}^{(t)}\leftarrow\{C_{j}\in\mathcal{C}_{n^{*}}^{(t)}\mid\mathbf{g}_{n^{*}}^{(t)}[j]=1\}$
13:     $\mathcal{C}\leftarrow\mathcal{C}\cup\mathcal{C}_{\text{good}}^{(t)}$
14:     $\mathcal{S}^{(t+1)}\leftarrow\mathcal{S}^{(t)}\setminus\bigcup_{C\in\mathcal{C}_{\text{good}}^{(t)}}C$
15:     $t\leftarrow t+1$
16:end while
17:return $(\mathcal{C})$
At iteration $t$, the current set of unassigned sentences is denoted as $\mathcal{S}^{(t)}$. A special case in the first iteration, where $\mathcal{S}^{(0)}=\mathcal{S}$ represents the entire set of unique sentences derived from the dialogue corpus. For each candidate cluster number $n_{i}\in N$, we compute:
$\mathcal{C}_{n_{i}}^{(t)}=F(\mathcal{E}^{(t)},n_{i}),$
where $F$ is a clustering function (e.g., K-means clustering), and $\mathcal{E}^{(t)}=\{\mathbf{e}_{s}\mid s\in\mathcal{S}^{(t)}\}$ representing the sentence embeddings. This results in $N$ distinct cluster assignments.
Then, the semantic coherence of each cluster is evaluated using a fine-tuned LLM $\mathcal{M}_{\text{eval}}$:
$\mathbf{g}_{n_{i}}^{(t)}=\left[\mathcal{M}_{\text{eval}}(C_{1}^{(t)}),\ldots,\mathcal{M}_{\text{eval}}(C_{n_{i}}^{(t)})\right],$
where $\mathcal{M}_{\text{eval}}(C_{k})=1$ if the cluster is coherent (i.e., “good” cluster), else 0.
While the number of clusters is often known with benchmark evaluation, it is challenging but necessary to implement mechanisms for determining the optimal number of clusters in a noisy text corpus. Here, we propose a local search heuristic that maximizes the “good/bad” ratio at each iteration. The optimal $n_{t}^{*}$ is selected as:
$n_{t}^{*}=\arg\max_{n_{i}\in N}\frac{\sum_{j=1}^{n_{i}}\mathbb{I}\left(\mathbf{g}_{n_{i}}^{(t)}[j]=1\right)}{\sum_{j=1}^{n_{i}}\mathbb{I}\left(\mathbf{g}_{n_{i}}^{(t)}[j]=0\right)+1},$
where $n_{t}$ denotes the best cluster number at iteration $t$. This enables the automatic discovery of suitable cluster numbers in a step-by-step manner. The search space $\bm{N}$ should be chosen carefully to balance accuracy and efficiency; thus, we propose a search space pruning strategy, as illustrated in Section 5.6 for more details.
Finally, the “good” clusters in the optimal $\mathcal{C}_{n^{*}}^{(t)}$ are retained, and the remaining sentences will be refined in the next (i.e., $t+1$) iteration, enabling the iterative discovery of high-quality clusters. The proposed method is summarized in Algorithm 1, and the LLM integration is highlighted.
3.3 Post-Correction with LLM-Generated Intent Labels
Preliminary results in Table 12 suggest that multiple clusters may capture similar intents due to diminishing size of unassigned sentences $\mathcal{S}^{(t)}$ in later iterations, leading to smaller and less representative clusters. The embedding distances between sentences within clusters limit the natural consolidation of clusters with similar intents but different expressions Khan et al. (2020), necessitating a post-correction step to merge semantically aligned clusters. In contrast to previous methods that rely on direct LLM queries for validating individual cluster assignments Viswanathan et al. (2024); Feng et al. (2024), which are computationally expensive and inconsistent, we propose a context-aware approach, leveraging LLMs’ naming utility to robustly merge clusters based on intent labels.
At the end of the iterative clustering process, each cluster $C_{k}$ receives an intent label from the fine-tuned LLM naming utility:
$l_{k}=\mathcal{M}_{\text{name}}(C_{k}).$
These labels concisely summarize the semantic information of each cluster and are mapped to the semantic space of sentence embeddings using the embedding function $f_{\text{emb}}$:
$\mathbf{l}_{k}=f_{\text{emb}}(l_{k})\in\mathbb{R}^{d},\quad\|\mathbf{l}_{k}\|_{2}=1.$
This normalization positions label embeddings $\mathbf{l}_{k}$ on the unit hypersphere $\mathbb{S}^{d-1}=\{\mathbf{x}\in\mathbb{R}^{d}\mid\|\mathbf{x}\|_{2}=1\}$, enabling accurate measurement of semantic relationships along the sphere’s surface rather than through straight-line distances Fletcher et al. (2004), thus facilitating precise intent similarity comparisons in high-dimensional space.
Then, the clusters are structured into a semantic affinity graph $G=(V,E)$ to model relationships based on label similarity. Vertices $V=\{C_{1},C_{2},\ldots,C_{K}\}$ represent clusters, and the edges $E\subseteq V\times V$ are determined using hyperspherical geometry by computing the geodesic distance, which captures angular separation between label embeddings Fletcher et al. (2004):
$\text{Dist}(\mathbf{l}_{i},\mathbf{l}_{j})=\arccos(\langle\mathbf{l}_{i},\mathbf{l}_{j}\rangle),$
forming edges if $\text{Dist}(\mathbf{l}_{i},\mathbf{l}_{j})<\theta$, with $\theta$ is a predefined threshold ($\theta=0.8$).
For each derived edge, a probabilistic criterion is designed to enhance the robustness of merging decisions, naturally modeling label embeddings as samples from a mixture of von Mises-Fisher distributions Banerjee et al. (2005):
$p(\mathbf{l}_{k}\mid\boldsymbol{\mu}_{m},\kappa_{m})=Z_{d}(\kappa_{m})\exp(\kappa_{m}\langle\mathbf{l}_{k},\boldsymbol{\mu}_{m}\rangle),$
where $\boldsymbol{\mu}_{m}=\mathbf{l}_{m}\in\mathbb{S}^{d-1}$ is the mean direction of the $m$-th intent embedding, $\kappa_{m}>0$ controls the distribution’s tightness, and the normalization constant $Z_{d}(\kappa_{m})$ ensures the density integrates to 1 over the hypersphere:
$Z_{d}(\kappa)=\frac{\kappa^{d/2-1}}{(2\pi)^{d/2}I_{d/2-1}(\kappa)}.$
Thus, the probability that clusters $C_{i}$ and $C_{j}$ share the same intent is computed for each edge as:
$P(\text{same}\mid\mathbf{l}_{i},\mathbf{l}_{j})=\sum_{m=1}^{K}\pi_{m}p(\mathbf{l}_{i}\mid\boldsymbol{\mu}_{m},\kappa_{m})p(\mathbf{l}_{j}\mid\boldsymbol{\mu}_{m},\kappa_{m}),$
with $\pi_{m}=1/K$ as uniform mixture weights. An edge is retained only if this probability exceeds a predefined threshold $\tau$ ($\tau=0.7$), minimizing inappropriate merges by accounting for uncertainty.
Finally, clusters linked in the affinity graph are consolidated into connected components, forming a refined cluster assignment $\mathcal{C}^{\prime}=\{C^{\prime}_{1},C^{\prime}_{2},\ldots,C^{\prime}_{K^{\prime}}\}$ that eliminates redundancy, enhances interpretability, and maintains robust semantic alignment. Each merged cluster $C^{\prime}_{k}$ receives a new intent label reflecting its semantic content.
3.4 Context-Aware Role Separation with LLM-Generated Intent Labels
From a practical perspective, dialogue datasets are often accompanied by domain-specific features or use-case scenarios (e.g., customer service calls, group discussions). This motivates the incorporation of contextual information into the clustering process. The customer service dialogues typically involve only the customer and the service agent Lin et al. (2022). The classification task that assigns a sentence to its associated role is relatively simple with labeled training data. To provide an unsupervised solution, the sentence roles are naturally determined based on the LLM- intent label. By heuristics, sentences within clusters labeled as “inquire-” or similar actions are mostly sent from the customer, and vice versa. A two-step approach is proposed: initially, intermediate clustering results are obtained using previous approach, denoted as $C_{\text{inter}}=\{C_{1},C_{2},\ldots,C_{K}\}$, and these results are divided based on intent labels $l_{k}$, forming two groups with distinct roles, $R_{\text{customer}}$ and $R_{\text{agent}}$, where:
$\begin{array}[]{c}R_{\text{customer}}=\{s\in S\mid l_{k}(s)\in\{\text{"inquire-"},\dots\}\},\\ R_{\text{agent}}=\{s\in S\mid l_{k}(s)\notin\{\text{"answer-"},\dots\}\}.\end{array}$
In the second step, the sentences corresponding to each role are clustered again, denoted as $C^{\prime}_{\text{customer}}=\text{Cluster}(R_{\text{customer}})$ and $C^{\prime}_{\text{agent}}=\text{Cluster}(R_{\text{agent}})$, to achieve more refined outcomes, allowing for a clear separation of customer intent from service agent intent. Finally, the two groups of clusters are merged together, serving as the final or intermediate result to be further processed.
4 Experiment
4.1 Dataset
The proposed dataset contains 1,507 high-quality intent clusters manually annotated from over 100,000 realistic customer service calls, comprising 55,085 distinct sentences with an average length of 17 Chinese characters per sentence (see Appendix A for annotation details) [^1] The full dataset and fine-tuned models will be released upon acceptance. Sample dataset can be found at GitHub.. Among these intents, 885 are identified as domain-specific queries (e.g., inquire-insurance), while 622 are categorized as out-of-domain (e.g., provide-location), representing general queries that commonly arise in a customer service interaction. This dataset is the first Chinese benchmark for intent clustering and the largest of its kind compared to the existing English benchmark datasets shown in Table 1. It presents several new challenges, including the lexical sparsity in short-text sentences, the dependence on contextual knowledge, and the difficulties in balancing accuracy with computational efficiency due to the excessively large intent size. The semantic diversity, calculated based on the average cosine distance between individual sentences and the cluster centroid Casanueva et al. (2022), depicts the complexity of this dataset and motivates the development of semantic-guided techniques.
Number of Number of Semantic
Dataset sentences intents diversity
BANK77 3080 77 0.209
OOS 23,700 150 0.254
NLU++ 3,080 62 0.367
CLINC(I) 4,500 150 0.275
MTOP(I) 4,386 102 0.234
Massive(I) 2,974 59 0.351
ours 55,085 1507 0.538
Table 1: Comparison of proposed dataset with existing intent clustering datasets.
Category Method NMI #Good
Baselines K-Means 0.7899 94.8%
GMMs 0.7903 91.1%
Hierarchical* 0.8001 94.9%
LLM-Guided Clustering Baselines ClusterLLM Zhang et al. (2023) 0.7284 91.2%
IDAS De Raedt et al. (2023) 0.8109 93.6%
Keyphrase Viswanathan et al. (2024) 0.8371 94.5%
LLMEdgeRefine Feng et al. (2024) 0.7411 87.2%
Proposed Method (context-free) LLM-ITL (random) 0.8202 97.7%
LLM-ITL (convex) 0.8208 97.8%
LLM-ITL + keyphrase 0.8378 96.4%
Proposed Method (context-aware) LLM-ITL + merge 0.8420 97.8%
LLM-ITL + role 0.8679 97.2%
LLM-ITL + role + merge 0.8826 97.6%
Table 2: Performance of baselines and proposed methods in dialogue intent clustering. The best results are bolded.
{CJK}
UTF8gkai
LLM qwen7b qwen14b baichuan2-7b chatglm3-6b
Accuracy 96.25% 97.50% 89.17% 95.83%
Table 3: Performance of fine-tuned LLMs in evaluating semantic coherence of intent clusters.
{CJK}
UTF8gkai
LLM qwen7b qwen14b baichuan2-7b chatglm3-6b
Accuracy 92.8% 94.3% 94.3% 94.4%
Table 4: Performance of fine-tuned LLMs in naming intent clusters.
4.2 Metrics
Normalized Mutual Information (NMI) measures the degree of similarity between ground-truth and predicted clusters, scaled between 0 (no mutual information) and 1 (perfect correlation). However, researchers have claimed biased behavior in NMI Jerdee et al. (2024), necessitating further evaluation that reflects the true quality of the intent cluster from human perception.
Goodness score measures the proportion of good clusters evaluated by the fine-tuned LLM evaluator. For intermediate steps of iteration where the number of clusters is unknown, the good/bad ratio is used as an invariant measure of clustering quality:
$\text{goodness}_{i}=\frac{\text{\# good clusters}}{\text{\# bad clusters}}$
For the final clustering results, the percentage of good clusters among all clusters is reported. Note that the evaluation uses a different LLM evaluator than the one used for intermediate evaluation to ensure the fairness of the reported metric:
$\text{goodness}_{final}=\frac{\text{\# good clusters}}{\text{\# {total clusters}}}$
This metric offers a comprehensive understanding of the quality of the produced intent clusters, allowing for accurate cluster-level evaluation. Additionally, it can be easily deployed in applications without ground-truth annotations, thereby eliminating the need for human involvement.
4.3 Implementations
Four open-sourced Chinese LLMs [^2] The models are available at: Qwen2.5-7B; Qwen2.5-14B; Baichuan2-7B ; ChatGLM3-6b are fine-tuned using LoRA Hu et al. (2022) on 4 $\times$ Nvidia A100 GPUs. For coherence evaluation, a human-annotated training dataset of 1,772 intent clusters labeled as “good” or “bad” is used. For cluster naming, a training dataset of 2,500 clusters, each containing 20 sentences, is annotated with intent labels. Semantic-rich embeddings are generated using the BAAI General Embeddings (BGE) model [^3] https://huggingface.co/BAAI/bge-large-zh-v1.5 Xiao et al. (2024). The best-performing clustering algorithm, hierarchical clustering, and LLM utilities, qwen14b and chatglm3-6b, are selected for LLM-in-the-loop clustering. To reduce computational load, both random and convex sampling methods are implemented to select 20 representative sentences per cluster as input for the LLM.
5 Results and Discussions
5.1 LLM Utilities
Table 3 presents the performance of the fine-tuned LLM coherence evaluator tested on 480 unseen clusters. The results indicate that mainstream open-source LLMs can effectively serve as robust evaluators for assessing the quality of intent clusters and providing human-aligned judgments. For cluster naming, since labels are not unique, accuracy is manually evaluated by four human experts based on alignment with the true labels in the dataset. The results in Table 4 demonstrate that the fine-tuned LLMs show promising performance in generating intuitive names and adhering strictly to the predefined “Action-Objective” format.
5.2 Main Results
Table 2 presents the main results evaluated using the proposed dataset. Among the LLM-guided clustering baselines, the data-centric Keyphrase expansion method, which refines sentence embeddings via LLM-summarized keyphrases, achieved the most significant performance gain Viswanathan et al. (2024). However, deriving pairwise constraints for fine-tuning the embedding model, as in ClusterLLM, was ineffective due to the excessive number of clusters in the dataset. Interestingly, goodness evaluation does not always align with the NMI score. For example, while keyphrase expansion improved NMI, the generated keyphrases could distort the original sentence meaning, leading to less representative clusters and a slightly lower goodness score compared to the base model. This highlights the need to balance ground truth alignment with semantic coherence during evaluation.
The proposed LLM-in-the-loop intent clustering demonstrated satisfactory improvements in NMI and significant enhancements in the goodness score. In a context-free setting, iterative clustering with convex sentence sampling outperformed both ClusterLLM and IDAS baselines. Effective keyphrase data augmentation further improved NMI, highlighting the potential of integrating data-centric methods into the pipeline. Furthermore, the context-aware approaches consistently improved performance by incorporating dialogue roles and cluster merging, resulting in a 4.48% increase in NMI without compromising the goodness score. This emphasizes the importance of task-centric design in leveraging intent labels effectively for the problem of intent clustering.
Notably, the proposed method does not require prior knowledge of the number of clusters and performs parameter searches during clustering, unlike baseline methods that rely on the true number of clusters. This adaptability underscores its strong application potential for extracting intent clusters from noisy text corpora Akama et al. (2020).
Method Bank77 CLINC(I) MTOP(I) Massive(I)
SCCL 81.77 92.94 73.52 73.90
Self-supervise 80.75 93.88 72.50 72.88
ClusterLLM 85.07 94.00 73.83 77.64
IDAS 82.84 92.35 72.31 75.74
LLMEdgeRefine - 94.86 72.92 76.66
ours 82.32 94.12 72.45 78.12
Table 5: Performance (NMI) on English benchmarks.
5.3 Evaluation on English Benchmarks
To validate the effectiveness of the proposed LLM-in-the-loop technique, we evaluate the context-aware (best-performing) method on four widely used English benchmarks FitzGerald et al. (2023). The Llama-7b model is fine-tuned with 800 intent clusters (200 samples from each dataset) annotated by human experts to derive Action-Objective intent labels and bad clusters through perturbation. As shown in Table 5, our method delivers performance comparable to state-of-the-art LLM-guided methods and outperforms on Massive(I) dataset. Furthermore, the computational cost is lower than the compared methods as measured in Section 5.5.
5.4 Application Performance
One objective of intent clustering is to create a high-quality labeled dataset for training intent classifiers that can effectively handle future data inputs Gung et al. (2023). To demonstrate the practical advantages of the proposed methods, we further evaluate their performance by training BERT classifiers (bert-base-chinese model) based on the clustered results. Compared to the training data produced by the base model, which achieved only 65% classification accuracy, the proposed LLM-in-the-loop method improved the accuracy by 12%, providing promising evidence of the improved cluster quality.
5.5 Analysis of Computational Cost
We systematically compare the cost of the proposed method with existing LLM-guided clustering on the Bank77 dataset, with a test set of $S=3,080$ sentences and true cluster number $N=77$. The proposed LLM-in-the-loop iterative method (Algorithm 1) uses candidate cluster numbers $N=[10,30,50,70]$, requiring $\sum_{n_{i}\in N}n_{i}=160$ LLM calls per iteration, and with $T=3$ iterations, totals $T\cdot\sum_{n_{i}\in N}n_{i}=480$ calls, each with 20 sentences input for coherence evaluation. ClusterLLM, employing a fixed triplet sampling approach with $Q=1,024$, incurs a constant $1,024$ LLM calls. Keyphrase expansion generates one keyphrase per sentence, resulting in a linear cost of $3,080$ LLM calls for the entire dataset. While our method requires more input tokens per LLM call, the cost remains efficient compared to its counterparts since the output token (i.e., “good/bad” label and intent label) is kept concise with fewer LLM calls.
5.6 Ablation Studies
Effective Coherence Evaluation with Data Sampling and LLM Crowdsourcing
Optimizing the input to LLM evaluator is crucial for enhancing performance and reducing costs Song et al. (2025); Jiang et al. (2024). As shown in Figure 3, convex sampling outperforms random sampling in selecting representative sentences for coherence evaluation. Interestingly, as the sample time $t$ for repeated evaluation with majority voting or the dimension of the convex hull $d$ increases, performance consistently declines. This suggests that simpler, more efficient parameters lead to better accuracy. This pattern is verified through experiments with the ChatGLM model (see red line), which exhibits the same pattern.
Figure 3: Performance comparison of sampling methods and hyperparameters for effective LLM evaluation.
To better assess the consistency and alignment between LLM judgments and human perceptions, we utilize the concept of crowdsourcing for collaborative coherence evaluation Zhang et al. (2025). The proposed LLM crowd consists of diverse entities Zhang et al. (2024) represented by four fine-tuned LLMs and uses a majority voting mechanism to aggregate their judgments Schoenegger et al. (2024). Based on the evaluation, we argue that a single fine-tuned LLM can effectively serve as a robust evaluator, as evidenced by the comparable performance to crowdsourced accuracy (cross markers in Figure 3) when using convex sampling.
LLM qwen7b qwen14b baichuan2-7b chatglm3-6b
Accuracy 62.50% 65.00% 58.50% 60.00%
Table 6: Performance of fine-tuned LLMs in evaluating semantic coherence of intent clusters using numerical 1–4 scale scoring.
Binary Judgment vs. Numerical Scoring
To identify the optimal evaluation strategy for our clustering framework, we compared a numerical 1–4 scoring system (1: very poor, 4: very coherent) with the proposed binary good/bad judgment. We annotated 1,000 clusters with input from five human experts, finalized scores through majority voting, and fine-tuned four Chinese LLMs using this dataset. The models were then evaluated on 200 additional clusters, assigning scores five times per cluster and consolidating results via majority voting. The 1–4 scale demonstrated low accuracy and inconsistency, with 19% of samples observed to have significant disagreement (e.g., three ‘1’s and two ‘4’s) in Qwen-2.5-14B, making it unreliable for assessing cluster quality or aligning with human evaluation. In contrast, the simpler binary good/bad judgment provided consistent and interpretable quality assessments, proving to be the more effective evaluation strategy.
Method NMI Goodness
Geodesic distance (probabilistic) 0.8420 97.8%
Cosine similarity (deterministic) 0.8102 94.7%
Cosine similarity (probabilistic) 0.8242 95.2%
Geodesic distance (deterministic) 0.8309 95.4%
Table 7: Ablation study on cluster merging methods.
Effectiveness of Cluster Merging Techniques
Table 7 evaluates the impact of distance measures and merging decisions on the proposed cluster merging technique. Results highlight the superiority of geodesic distance in hyperspherical space, demonstrating its ability to effectively capture deeper semantic relationships compared to cosine similarity. Additionally, the probabilistic merging strategy consistently improves robustness by accounting for uncertainty, yielding better clustering quality regardless of the distance measure used. These findings emphasize the importance of both an appropriate distance metric and a probabilistic approach in achieving optimal merging outcomes.
{CJK}
UTF8gkai
Top1 Top2 Top3 Top4 Top5
Accuracy 26.32% 47.37% 73.68% 84.21% 89.47%
Table 8: Search space pruning by LLMs.
Parameter Search Space Pruning
In practice, the search space for the number of clusters can be extensive, ranging from a few options to several hundred. To mitigate the computational costs associated with traditional model-based selection, we propose an LLM-native method for search space pruning, which predicts the optimal cluster number for subsequent iterations using logs from previous iterations (see Appendix 13). By pruning the search space and selecting the top 5 non-repetitive predicted solutions, we achieved an accuracy of 89.47% (see Table 8), indicating a high likelihood that the optimal cluster number is among predicted candidates. This approach substantially reduces redundant model fittings and enhances the efficiency of deploying the proposed methods.
6 Conclusion
This paper tackles dialogue intent clustering using a human-aligned LLM-in-the-loop approach. Experiments on a complex, realistic Chinese customer service dataset reveal that fine-tuned LLM utilities excel in semantic coherence evaluation and cluster labeling, with the proposed methods consistently outperforming existing LLM-guided clustering baselines in cluster quality and computational efficiency. Future work should focus on refining evaluation strategies to enhance semantic guidance from LLMs and exploring task-centric approaches for deeper LLM-in-the-loop integration.
Limitations
This work has two main limitations. First, while the coherence evaluation is practical and intuitive, relying solely on cluster coherence as the quality indicator overlooks other critical attributes, such as meaningfulness and expressiveness, which are equally important in assessing a cluster’s usefulness. For instance, a cluster labeled “express-feeling” may be overly general and should be refined into more specific categories like “express-appreciation” and “state-dissatisfaction” to enhance interpretability and practical applicability. Additionally, the binary evaluation and numerical scoring considered in this paper are deterministic indicators, overlooking probabilistic judgments and confidence intervals, which could provide deeper insights into clustering assessment and offer greater flexibility for integration into the LLM-in-the-loop clustering process.
Second, the dataset proposed in this study is limited to the Chinese language and focuses on specific domains, namely banking, insurance, and telecommunication. While it represents the largest realistic customer service call dataset for the intent clustering problem, its coverage is restricted in terms of multi-domain and multilingual aspects, potentially affecting the generalizability of our findings to other languages and domains. Furthermore, the scope of this paper is confined to intent clustering, which serves as the initial step in the broader process of intent discovery. Subsequent steps, such as analyzing intent trajectories and recognizing intents at the document level rather than the sentence level, are not comprehensively discussed. These areas warrant further investigation to fully reveal the practical values of LLM-in-the-loop design in intent mining tasks.
Appendix A Data Collection and Annotation
This section details the three-stage data collection and annotation process for the proposed Chinese dialogue intent clustering dataset.
Data Collection.
The original dialogue dataset is sourced from audio transcriptions of customer service calls across several domains. Initially, the dataset contains over 11,879 customer service calls. Strict filtering is applied to maintain data integrity and exclude confidential or sensitive content, resulting in 8,184 dialogues and 69,839 sentences. Subsequent data cleaning removes duplicate sentences, resulting in a final set of 55,085 unique sentences.
Data Annotation.
To construct the intent clustering dataset, we internally recruited 15 human experts, each with over five years of experience in customer service call centers, to participate in the data annotation process:
  1. 1.
    K-means clustering was initially applied with $n=2000$, serving as a starting point for the data annotation.
  2. 2.
    Each expert assessed the initial clusters by inspecting the semantic coherence of the sentences. They were instructed: “Label the cluster as ’Good’ if all sentences share the same intent; otherwise, label it as ’Bad’. This resulted in 1283 good clusters and 717 bad clusters.
  3. 3.
    For the good clusters, annotators were asked to label the underlying intent using the naming convention “Action-Objective.” Five samples were provided as guides. After intent annotation, 1255 clusters with unique intentions were retained, and 28 clusters were merged due to replicated intent.
  4. 4.
    For the bad clusters, annotators reassigned each sentence to the appropriate intent cluster based on the annotated labels. Sentences that did not fit into any preexisting clusters were assigned to new clusters, and the same process was repeated as in Step 2.
Data Verification.
In the final stage, a separate group of 10 experts reviewed the annotated clusters for accuracy and consistency. Any discrepancies were resolved through consensus, ensuring the dataset’s reliability and validity for further analysis. The final dataset consists of 1,507 high-quality intent clusters.
Appendix B Prompt Template
The prompts for coherence evaluation and cluster naming are translated into English for demonstration purposes. Each input is accompanied by a few-shot demonstration with five input-output pairs to ensure consistency in the output format and enhance understanding of the task.
Coherence Evaluation -  Your are a helpful assistant for sentence clustering. Based on the relevancy and common points of the following sentences in a cluster, classify the cluster as: “Good” or “Bad”. Only provide the label without any additional content. Example: input:[sentences] output:[label] input:{[sentences]} output:
Cluster Naming -  Your are a helpful assistant for sentence clustering. Based on the relevancy and common points of the following sentences in a cluster, summarize the cluster with an “Action-Objective” label. Only provide the label without any additional content. Example: input:[sentences] output:[label] input:{[sentences]} output:
Appendix C More Results: Evaluation with Proprietary LLM
Given the resource-intensive nature of LLM fine-tuning, we conducted further experiments in coherence evaluation using proprietary LLMs accessed through OpenAI APIs: GPT-3.5, GPT-4, and GPT-4o. These advanced and widely used models offer strong performance in various tasks but also come with higher operational costs due to token consumption. Using the same evaluation set of 480 clusters, these models achieved accuracies of 89.58%, 93.54%, and 94.17%, respectively. The results highlight that a smaller, fine-tuned LLM, such as Qwen-2.5-7b (with an accuracy of 96.25%), can outperform these advanced models while significantly reducing costs associated with API usage [^4] The fine-tuned Chinese LLMs for coherence evaluation and cluster naming will be released upon acceptance.. For comparison, a vanilla Qwen-2.5-7b model (without fine-tuning) was also evaluated, yielding a substantially lower accuracy of 75.63%, emphasizing the essence of fine-tuning.
{CJK}
UTF8gkai
Cluster Coherence Original Sentences English Translation
Good "给企业固定资产买保险,大约能投大约得投保多少呢", "就是假如我有一百万的企业固定资金买保险大概投保多少钱啊", "请问一下我想咨>询一下企业财产保险", "您好我想了解一下这个企业财产保险", "就了解一下这个企业财产保险","你们是财险", "是哦这个属于财险了对吧", "企业财产保险的", "这些都属于财产险对吗", "财险人工那你这是", 就是如果我要为我这个私营企业买这个保险需要什么手续", "财产险", "呃企业财产保险是以什么为保" "If I buy insurance for a company’s fixed assets, how much insurance will it cost?", "That is, if I have a million corporate fixed assets, how much will it cost to buy insurance?", "Excuse me, I would like to consult> Ask about corporate property insurance", "Hello, I want to know about this corporate property insurance", "Just want to know about this corporate property insurance", "You are a property insurance company", "Yes, this belongs to property insurance "Right?", "Enterprise property insurance", "These all belong to property insurance, right?", "What about property insurance workers?", What do I need if I want to buy this insurance for my private enterprise? Procedure","Property Insurance", "Well, what does corporate property insurance cover?"
Bad "就是连连续,就是一直一直保", "但是它连不上是怎么回事啊?", "哦就是主要是直接给公司转账对吧?", "接吗", "接也是吗?", "我直>接去", "你是直接直接用那个" , "就是从哪接过来再接回去","还是需要从哪儿连线这个宽带", "你直接给我说这些啊", "直接把", "哦直接", "嗯那个礼品是直接就发放了", "是你直接给我回复对吗", "直接到那里去", "嗯,最好直飞。", "请帮我连接+", "把钱打过去的话,我是直接打到那个证券公司", "直接就是" "It’s continuous, it’s always guaranteed", "But what’s wrong with it not being able to connect?", "Oh, it’s mainly to transfer money directly to the company, right?", "Yes", "Yes too?" ?", "I’ll go directly", "Are you using that directly" , "Just connect it from where you are and then connect it back", "Or do you need to connect to the broadband from somewhere", "You Just tell me this directly", "Just give it directly", "Oh directly", "Well, the gift was given out directly", "You replied to me directly, right?","Go there directly" , "Well, it’s best to fly directly.", "Please help me connect +", "If I call the money, I will call the securities company directly", "Directly"
Table 9: Example of “Good” and “Bad” intent clusters in coherence evaluation.
{CJK}
UTF8gkai
Cluster Name Original Sentences English Translation
询问-优惠 (Inquire-Promotion) ’那有什么优惠券什么之类的吗?’, ’是是怎么形式是优惠吗?’, ’还是不是优惠活动,是那个直接给我打我卡里吗 ,还是是什么优惠券儿啊?’, ’就比如新用户他有什么优惠券儿之类的吧!’, ’嗯你你那还有什么优惠的活动吗? 就是比较合适就是合适的动。’ ’Are there any coupons or anything like that? ’, ’What is the form of the discount? ’, ’Still, it’s not a promotion. Is it the one that directly charges my card? , or is it some kind of coupon? ’, ’For example, what coupons does a new user have? ’, ’Well, do you have any other discounts? It is more appropriate and appropriate to move. ’
解答-金额 (Answer-Amount) ’一共是三十一块二毛’, ’就每个月一百三十八’, ’对,一个月也就是四百百四五百块钱嘛, 给您自己做个积累。’, ’十二月份的话是用了三十三块九毛二’, ’对一个月一百三十八’ ’The total is thirty-one and twenty cents’, ’That’s one hundred and thirty-eight cents per month’, ’Yes, that’s four hundred, four hundred and five hundred yuan per month. Make an accumulation for yourself. ’, ’In December, it cost thirty-three dollars and ninety-two cents’, ’That’s one hundred and thirty-eight dollars a month’
Table 10: Example of cluster naming with "Action-Objective" style.
Dataset Naming Convention Example (sentence) Example (label)
NLU scenario-intent Send an email to Alex and write thank you. email sendemail
NLU++ list of keywords How long does it usually take to get a new pin? ["how_long","pin", "arrival","new"]
OOS objective Please tell me why my card was declined yesterday. card_declined
ours action-objective Well, do you have any other discounts? inquire-promotion
Table 11: Comparison of cluster naming convention in intent clustering datasets.
{CJK}
UTF8gkai
Cluster Name Clusters with Similar Intention English Translation
询问-意外事故 (Inquire-Accident) "假如被车碰了或者是被楼上的砖砸了一下", "给别人儿撞的意外", "啊撞到别人然后就是", "就平时有时候开车嘛可能会遇到这个", "嗯哦这种情况,那要是就是我自己不小心撞到了那个某个地方然后", "就是把其他的东西撞到了呀什么的", "那人生意,不是就是,如果是不小心在马上被车撞了的话", "撞到人了吧", "然后不小在行驶当中被别人损害就是说拿石头砸的呀然后", "什么被车撞了之类的,是吗", "被被被撞了,还是被什么一些什么意外事故了", "我把别人的车撞了是吧" "If I were hit by a car or hit by a brick from upstairs," "Accidentally hit by someone else," "Ah, hit someone else and then," "Just sometimes when driving, you might encounter this," "Well, if I’m not careful and hit some place myself," "Just hit something else or something," "That’s a human affair, not just, if it’s accidentally hit by a car on the road," "Hit someone, right?" "Then not small in the process of driving, being damaged by someone else, say, hit with a stone, and then," "What, hit by a car or something, right?" "I hit someone else’s car, right?"
询问-意外死亡 (Inquire-Accident Death) "哦猝死,那猝死算意外吗", "啊,那我想知道那个猝死的话,算是意外死亡吗", "算意外死亡吗", "那如果猝死算是意外死亡吗?", "猝死也算意外事吧", "那那猝死是意外死亡吗如果是猝死的话是", "猝死算是意外死亡吗", "猝死算意外死亡吗", "嗯,那个猝死,猝死属于意外意外险吗,意外死亡吗", "那猝死的话,算意外死亡吗?", "那那那那个就是那个猝死算是意外死亡吗" "Oh sudden death, is sudden death considered accidental?" "Ah, I want to know if sudden death is considered accidental death?" "Is it considered accidental death?" "If sudden death is considered accidental death?" "Sudden death is also an accident, right?" "Is sudden death considered accidental death if it is sudden death?" "Is sudden death considered accidental death?" "Is sudden death considered accidental death?" "Well, that sudden death, does sudden death fall under accidental insurance, accidental death?" "Is sudden death considered accidental death?" "Is that, that, that sudden death considered accidental death?"
Table 12: Example of two high-quality intent clusters with similar intentions.
Epoch 1th 2th 3th
n_cluster good bad rate n_cluster good bad rate n_cluster good bad rate
20 20548 34537 0.595 20 0 2901 0.0 20 9 438 0.021
50 24100 30985 0.778 50 372 2529 0.147 50 70 377 0.186
100 32292 22793 1.417 100 804 2097 0.383 100 79 368 0.215
Best 1600 52184 2901 17.988 800 2454 447 5.490 200 86 361 0.238
Table 13: Example of log record for iterative intent clustering.